OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth, if necessary. x2 + 17 = 0

OpenStudy (austinl):

\[x^2 + 17 = 0\] \[x^2 = -17\] \[\sqrt{x^2} = \sqrt{-17}\] \[x = \sqrt{-17}\] \[x = i\sqrt{17}\]

OpenStudy (anonymous):

\[x^2 + 17 = 0 \rightarrow x^2 =- 17 \rightarrow x =\pm \sqrt{-17} \] i.e. \[x =\pm \sqrt{-1 \times17} \rightarrow x =\pm \sqrt{-1} \times \sqrt{17}\] \[x =\pm i \sqrt{17}\]

OpenStudy (anonymous):

how do I smipfy that because none of that is my answer choice

OpenStudy (uri):

No solution

OpenStudy (austinl):

Did you copy the question correctly?

OpenStudy (uri):

Do you have that choice?

OpenStudy (anonymous):

yes I did